i90i91+i92i93 note real numbers are exponents that add up to i to the 366 power
Answer by venugopalramana(3286)

i90i91+i92i93 note real numbers are exponents that add up to i to the 366 power
HOPE YOU MEAN
[I^90][I^91]+[I^92][I^93].ASSUMING SO
= [I^(90+91)]+[I^(92+93)]
=I^181+I^185
=I^(4*45+1)+I^(4*46+1)
=[(I^4)^45]I+[(I^4)^46]I...............SINCE I^4=[I^2]^2 = (-1)^2)=1,WE GET
=1*I+1*I=2I